$SPAD/src/input mset.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool mset.output
)set message test on
)set message auto off
)clear all
--S 1 of 17
macro I == Integer
--R
--R                                                                   Type: Void
--E 1
--S 2 of 17
macro symdif == symmetricDifference
--R
--R                                                                   Type: Void
--E 2
--S 3 of 17
s:Multiset I
--R
--R                                                                   Type: Void
--E 3
--S 4 of 17
t:Multiset I
--R
--R                                                                   Type: Void
--E 4
--S 5 of 17
t1:Multiset I
--R
--R                                                                   Type: Void
--E 5
--S 6 of 17
s := multiset [1,2,3,4,5,4,3,2,3,4,5,6,7,4,10]
--R
--R
--R   (6)  {1,2: 2,3: 3,4: 4,2: 5,6,7,10}

--R                                                      Type: Multiset(Integer)
--E 6
--S 7 of 17
t := multiset [2,2,2,9]
--R
--R
--R   (7)  {3: 2,9}
--R                                                      Type: Multiset(Integer)
--E 7
--S 8 of 17
union(s,t)
--R
--R
--R   (8)  {1,5: 2,3: 3,4: 4,2: 5,6,7,9,10}
--R                                                      Type: Multiset(Integer)
--E 8
--S 9 of 17
union(s,s)
--R
--R
--R   (9)  {2: 1,4: 2,6: 3,8: 4,4: 5,2: 6,2: 7,2: 10}
--R                                                      Type: Multiset(Integer)
--E 9
--S 10 of 17
intersect(s,t)
--R
--R
--R   (10)  {5: 2}
--R                                                      Type: Multiset(Integer)
--E 10
--S 11 of 17
difference(s,t)
--R
--R
--R   (11)  {1,3: 3,4: 4,2: 5,6,7,10}
--R                                                      Type: Multiset(Integer)
--E 11
--S 12 of 17
symdif(s,t)
--R
--R
--R   (12)  {1,3: 3,4: 4,2: 5,6,7,9,10}
--R                                                      Type: Multiset(Integer)
--E 12

--S 13 of 17
symdif(s,s)
--R
--R
--R   (13)  {}
--R                                                      Type: Multiset(Integer)
--E 13
--S 14 of 17
t1 := multiset [2,2]
--R
--R
--R   (14)  {2: 2}
--R                                                      Type: Multiset(Integer)
--E 14
--S 15 of 17
[t1 < t, t1 < s, t1 <= t, t1 <= s]
--R
--R
--R   (15)  [true,true,true,true]
--R                                                          Type: List(Boolean)
--E 15
--S 16 of 17
t1 := multiset [2,2,2]
--R
--R
--R   (16)  {3: 2}
--R                                                      Type: Multiset(Integer)
--E 16
--S 17 of 17
[t1 < t, t1 < s, t1 <= t, t1 <= s]
--R
--R
--R   (17)  [true,false,true,true]
--R                                                          Type: List(Boolean)
--E 17
)spool
)lisp (bye)

———-

References^{[1] nothing}

6